Helium filled balloon rises in the air because the density of helium is less than the density of air.

a. If air has an average molar of 29.0g/mol, what is the density of air at 25 and 770 mmHg?


1
Expert's answer
2021-01-05T05:57:43-0500

for ideal gas , we have

PV = n RT

here P is the pressure of gas , V is the volume of gas , n is the number of moles of gas taken.

by putting n = 1, and by putting the values of pressure and temperature , we can find the volume of 1 mole of gas .

mass of one mole of gas is equal to the molar mass of gas . ;ets it be represented by M .

Here,

P = 770mm Hg

T= (25+273)K = 298 K

n=1

R= 62.36 L mm Hg/mol . K


putting the values in ideal gas equation ,

770 V = 1 (62.36 ) 298

V = 24.13 L


To solve this question , we are assuming ideal gas behaviour and so ideal gas equation is applicable for all the gases given here .


meaning , one mole of every gas (air,helium,carbon dioxide) will occupy 24.13L at the given conditions


a). mass of one mole gas = 29g


density of gas = "\frac{Mass}{Volume}" = "\frac{29g}{24.13L}"


= 1.2 g/L